OpenStudy (anonymous):

Differentiate f(x)=(6-5x-x^2)/(x^2-1)

OpenStudy (psymon):

Well, we can factor both the top and bottom: 6 - 5x - x^2 = -x^2 - 5x + 6 = -(x^2+5x - 6) This then factors into -(x+6)(x-1) Now the bottom is a difference of squares and simply factors into (x+1)(x-1), meanign we have: \[\frac{ -(x+6)(x-1) }{ (x+1)(x-1) }= \frac{ -(x+6) }{ (x+1) }\] Now we still have to do quotient rule, but this makes the differentiation simpler. Do you know the formula for the quotient rule?

OpenStudy (anonymous):

I do, thanks!

OpenStudy (psymon):

Okay, cool. Try it out and see what ya get :3